Presented below is information related to Tamarisk Company. Date Ending Inventory (End-of-Year Prices) Price Index December 31, 2017 $ 87,400 100 December 31, 2018 212,625 189 December 31, 2019 209,304 216 December 31, 2020 235,170 234 December 31, 2021 281,736 252 December 31, 2022 339,039 261 Compute the ending inventory for Tamarisk Company for 2017 through 2022 using the dollar-value LIFO method. Ending Inventory 2017 $enter a dollar amount 87400 2018 $enter a dollar amount 134839 2019 $enter a dollar amount 105355 2020 $enter a dollar amount 118054 2021 $enter a dollar amount 111800 2022 $enter a dollar amount 129900
Answer #1

calculate following

Ending inventory
2017 87400
2018 212625/1.89 = 112500 87400*1+25100*1.89 = 134839
2019 209304/2.16 = 96900 87400*1+9500*1.89 = 105355
2020 235170/2.34 = 100500 113779
2021 281736/2.52 = 111800 142255
2022 339039/2.61 = 129900 189496